OpenStudy (anonymous):

How to prove identity: sin x-1/sin x+1 = -cos^2 x / (sin x+1)^2?

OpenStudy (anonymous):

\[\sin x-1divsin x+1 = -\cos ^{2}x \div(\sin x+1)^{?}\] \[\sin x-1/\sin x+1 = -1-\sin ^{2}x/(sinx+1)(sinx+1)\]

OpenStudy (anonymous):

that's all I've got up to and I'm stuck D:

OpenStudy (anonymous):

\[\LARGE {\sin x-1\over \sin x+1}={-\cos^2x\over (\sin x+1)^2}\] is this the given? :)

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

ok... on the right side you have the numerator \(-\cos^2 x\) from identity \(\sin^2x+\cos^2x=1\) you have: \[\LARGE \cos^2x=1-\sin^2x\] so in your case you have: \[\LARGE -1\cdot \cos^2x\] \[\LARGE -1\cdot (1-\sin^2x)\] so that becomes : \[\LARGE \frac{\sin x-1}{\sin x+1}=\frac{-1(1-\sin^2x)}{(\sin x+1)^2}\] are we good till here :)

OpenStudy (anonymous):

noo :( Do i switch the spots -1(1-sin^2x)/(sinx+1)^2 to (sinx+1)^2/-1(1-sinx^2) .. oh I'm confused D:

OpenStudy (anonymous):

no you don't .... from this: \[\large \frac{{\sin x - 1}}{{\sin x + 1}} = \frac{{ -\cos^2x }}{{{{\left( {\sin x + 1} \right)}^2}}}\] we've came up with... \[\large \frac{{\sin x - 1}}{{\sin x + 1}} = \frac{{ - 1 \cdot \left( {1 - {{\sin }^2}x} \right)}}{{{{\left( {\sin x + 1} \right)}^2}}}\]

OpenStudy (anonymous):

have you ever heard about this identity \[\LARGE \sin^2x+\cos^2x=1\] ?

OpenStudy (anonymous):

yes I have

OpenStudy (anonymous):

do you understand how this came up? \[\LARGE \frac{{\sin x - 1}}{{\sin x + 1}} = \frac{{ - 1 \cdot \left( {1 - {{\sin }^2}x} \right)}}{{{{\left( {\sin x + 1} \right)}^2}}}\]

OpenStudy (anonymous):

@milliex51

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I understand

OpenStudy (anonymous):

ok... nice. Now we multiply the numerator of right side which from: \[\LARGE \frac{{\sin x - 1}}{{\sin x + 1}} = \frac{{ - 1 \cdot \left( {1 - {{\sin }^2}x} \right)}}{{{{\left( {\sin x + 1} \right)}^2}}}\] becomes: \[\LARGE \frac{{\sin x - 1}}{{\sin x + 1}} = \frac{{{{\sin }^2}x - 1}}{{{{\left( {\sin x + 1} \right)}^2}}}\] are we good till here? :)

OpenStudy (anonymous):

um, bah. I guess so, do I: cancel sin^2x-1 from one of sinx+1? O.O

OpenStudy (anonymous):

don't hurry... now we can solve this in two ways... one of them is using cross multiply ... \[\LARGE \frac{{\sin x - 1}}{{\sin x + 1}} = \frac{{{{\sin }^2}x - 1}}{{{{\left( {\sin x + 1} \right)}^2}}}\] we get: \[\large \left( {{{\sin }^2}x - 1} \right)\left( {\sin x + 1} \right) = \left( {\sin x - 1} \right){\left( {\sin x + 1} \right)^2}\] \[\large \left( {{{\sin }^2}x - 1} \right)\cancel{\left( {\sin x + 1} \right)} = \left( {\sin x - 1} \right){\left( {\sin x + 1} \right)^{\cancel{2}}}\] \[\large \left( {{{\sin }^2}x - 1} \right) = \left( {\sin x - 1} \right){\left( {\sin x + 1} \right) }\] and now multiplying the right side by each other... we get: \[\left( {{{\sin }^2}x - 1} \right) = \left( {{{\sin }^2}x - 1} \right)\] the second way... \[\LARGE \frac{{\sin x - 1}}{{\sin x + 1}} = \frac{{{{\sin }^2}x - 1}}{{{{\left( {\sin x + 1} \right)}^2}}}\] \[\LARGE \frac{{\sin x - 1}}{{\sin x + 1}} = \frac{{{{\sin }^2}x - 1^2 }}{{{{\left( {\sin x + 1} \right)}^2}}}\] \[\LARGE \frac{{\sin x - 1}}{{\sin x + 1}} = \frac{{{{(\sin x+1)(\sin x-1 )}}}}{{{{\left( {\sin x + 1} \right)}^2}}}\] \[\LARGE \frac{{\sin x - 1}}{{\sin x + 1}} = \frac{{{{\cancel{(\sin x+1)}(\sin x-1 )}}}}{{{{\left( {\sin x + 1} \right)}^{\cancel{2}}}}}\] ...

OpenStudy (anonymous):

why did you square -1 from sin^2x-1?

OpenStudy (anonymous):

to show you that we can use this formula: \[\LARGE a^2-b^2=(a+b)(a-b)\] but you don't have to square it... :)

OpenStudy (anonymous):

really? ohh, thank you very much!

OpenStudy (anonymous):

My pleasure xD